# ChipletPart: Cost-Aware Partitioning for 2.5D Systems

### v2

@@ 32  Apply mutation to each offspring $\omega^{o}_{x}$ with mutation probability $p_{m}$;
33  Construct the next generation by selecting the top $\sigma$ elite genomes from the current generation, along with $k_{pop}$ offsprings;
34
@@ 4. Chiplet Partitioning Drivers
Smaller chiplets can potentially yield lower costs and improved yield, while technology heterogeneity offers better power and performance. The overhead of inter-chiplet IO also affects how a system should be partitioned into chiplets. In this section, we briefly discuss these factors.
@@ 4.1. Chiplet Cost Model
We use the open-source cost model from (Graening_CATCH, ; costmodelUCLA, ). It computes cost and yield based on the chiplet/interposer area (dependent on the technology node), assembly processes (dependent on bonding parameters), and IO placement constraints (dictated by netlist connectivity and reach). The model calculates the individual cost and yield of each chiplet along with the interposer, and then aggregates them for the full assembly cost. We consider designs as high volume in this case, to minimize the impact of NRE. The cost model is summarized by Equation 1. For further details, see (Graening_CATCH, ). Benefits of using the cost model in (Graening_CATCH, ) are discussed in Section 6.2 with results in Table 5.
@@ 4.2. Chiplet Power Model
We assume that performance of chiplets is preserved across technology nodes but that the power changes. We further assume that inter-chiplet communication latency is small enough to not influence the architecture. If certain inter-block interfaces are highly latency-sensitive, then those blocks should be merged in our framework so as to map them to the same chiplet. More complex performance models as in (ChenLZZL23, ) are possible but not explored in this work.
Power is calculated as the sum of block power and IO power. If all blocks are in the same chiplet, there will be no additional power due to IO. However, if the blocks are placed into multiple partitions (chiplets), some additional power will be consumed by the added IO cells. We scale the block power according to the scaling factors in (Stilmaker2017, ) for different technology nodes.
4.3. IO Reach Model
We use an IO cell model where each IO cell has both an area and a “reach” which is the maximum wire length that can be driven by the IO cell. The reach depends on the IO transceiver design and is often dictated by the inter-chiplet IO standard. E.g., Universal Chiplet Interconnect Express (UCIe) (UCIe1p0, ) guarantees a UCIe reach of $2\text{mm}$ for advanced packages while standard packages (such as organic substrates) support larger reaches of up to $25\text{mm}$.
Since we assume a fixed system architecture, our partitioner does not modify the network structure or insert pipeline stages to compensate for long inter-chiplet wires. We also assume a standardized IO cell with a fixed maximum reach instead of more complicated IO structures. These assumptions help keep the partitioning problem well-scoped and tractable. We leave sophisticated IO planning including pass-through connections, inserting buffers on the substrate, multiple IO types per design, etc. to future work. However, users can explore multiple architectural variants and IO types for the same design. For example, in our industry testcase (Section 6.1), we evaluate two versions that differ only in their crossbar configurations — and in Figure 13, we examine the impact of different IO cell types on partitioning.
@@ 4.3. IO Reach Model
[…]
Consider the example in Figure 4, and assume that a bundle of wires is connected from the side of one chiplet to the side of another. If the chiplets are $5\text{mm}$ on a side and spaced apart by $1\text{mm}$, then the short black connections require a reach of at least $1\text{mm}$, while the orange/blue/green connections respectively require reach of at least $7\text{mm}$$13\text{mm}$$19\text{mm}$. Small values of reach constrain the floorplan, while larger values of reach come at the cost of more expensive IO cells.
IOs are discretized depending on the standard. For instance, a single UCIe IO cell is as large as 0.88 $mm^{2}$ (UCIe1p0, ) while for parallel signaling (e.g., AIB or (Pal2021, )), an IO cell can be as small as 157 $\mu m^{2}$ (PalIO2021, ). This discretization can cause complexity for chiplet partitioning, as the IO (i.e., net cut) cost calculation follows a staircase-like curve as a function of net cut.
@@ 5. Chiplet Partitioning Core
In this section, we first discuss our chiplet-cost-driven partitioning approach ChipletPart-Core in Section 5.1. Then, we present our reach-aware chiplet floorplanning approach in Section 5.2.
@@ 5.1. Cost-driven Partitioning: Core-ChipletPart
Our Core-ChipletPart does not adopt the widely-used multilevel approach since there are typically only a few hundreds of IP blocks in a block-level netlist. As shown in Figure 5, we (i) compute a large pool of initial partitioning solutions using a variety of graph partitioning algorithms; (ii) prune poor-quality partitioning solutions from the pool; and (iii) perform floorplan-aware Fiduccia–Mattheyses (1585498, ) (FM)-based and Kernighan-Lin (HELSGAUN2000106, ) (KL) refinement. We then output the solution with the best cost.
Figure 5. Core-ChipletPart partitioning flow. We use multiple techniques shown in the partitioning pool to generate the initial partitions, then we prune out the worst-performing initializations before running refinement.
[…]
Generation of a partitioning pool. Similar to the initial partitioning stage in widely used multilevel partitioning approaches (KarypisAKS99, ) (BustanyGKKPW23, ), we compute a pool of initial partitioning solutions so as to explore a larger solution space. [^8] However, unlike multilevel approaches (KarypisAKS99, ) (BustanyGKKPW23, ), Core-ChipletPart omits the coarsening stage due to the relatively small number of vertices (i.e., 384 IP blocks in our largest testcase, as shown in Table 4). Specifically, we use the following graph partitioning methods.
  • Spectral partitioning: Because the block-level netlist is both small and sparse, we apply spectral partitioning (von2007tutorial, ). We first compute the spectral embedding using the two smallest nontrivial eigenvectors, then cluster the IP blocks with the K-means algorithm (likas2003global, ).999 In our implementation, we use a parallel K-means method with smart initialization (K-means++ (bahmani2012scalable, )), and we set $K=4$ by default.
  • High degree nodes expansion: Our node expansion approach follows (zhang2017graph, ). In the block-level netlist, high-degree nodes (e.g., crossbars) are distributed across different chiplets, and then expanded via breadth-first search. If multiple chiplets qualify for expansion, the block is assigned to the chiplet with which it has the strongest connection.
@@
Partitioning refinement. We use two different refinement strategies for refinement: (i) FM and (ii) KL. Both these processes are shown in Figure 6. Each potential FM move or KL pair swap triggers a call to the chiplet cost model to obtain the cost (gain) of an updated partitioning solution. This evaluation includes generating a floorplan solution using the “fast” mode of our SA-based reach-aware chiplet floorplanner, followed by calculating the corresponding cost. After completing each FM or KL pass, the “standard” mode of the SA-based reach-aware chiplet floorplanner is employed to search for an improved floorplan solution for the subsequent pass. Further details about our SA-based reach-aware chiplet floorplanner are provided in Section 5.2. We adopt the K-way FM implementation from TritonPart (BustanyGKKPW23, ), while our KL refinement (klrefine, ) follows the methodology described in (HELSGAUN2000106, ). In our implementation, we apply KL after FM to further improve solution quality, as KL explores a broader neighborhood by enabling pairwise vertex swaps — compared to FM’s single-vertex moves.
Figure 6. Floorplan-aware FM and KL-based refinement. In our implementation, we first run FM followed by KL.
[…]
@@ 5.2. Reach-aware Chiplet Floorplanning
Chiplet floorplanning is a crucial step, as it determines the location and shape of each chiplet on the interposer. In contrast to the well-studied macro placement, chiplet floorplanning must also consider IO feasibility. More specifically, the wirelength of a net must not exceed the reach specified by the corresponding IO cell (Section 4). Figure 7 illustrates the wirelength (measured in terms of Manhattan distance) calculation for net $e$ connecting chiplets $A$ and $B$. Assuming that the bitwidth of net $e$ is $n_{e}$ and the area of IO cells is $A_{IO}$, then the wirelength of net $e$ ($length(e)$) in the figure is calculated as: (i) $l=\sqrt{w^{2}+2A_{IO}}-w$ and (ii) $length(e)=h+2l$, where $l$ is the depth needed for all the IO cells. [^11] The detailed code for more general cases is available in (reach, ). Then the reach violation penalty for net $e$ is $n_{e}\times max(length(e)-reach(e),0.0)$. The reach violation penalty for a chiplet is defined as the summation of the reach violation penalties for all the nets connected to the chiplet.
Figure 7. Illustration of wirelength computation ($length(e)$). The yellow region indicates the placement area for IO cells that will satisfy the reach constraint.
[…]
@@
where $WL_{reach}$ represents the reach violation penalty, $A_{C}$ represents the area of chiplets, and $A_{P}$ represents the area of the package. $\alpha$, $\beta$ and $\gamma$ are user-defined coefficients that can be modified to achieve a desired trade-off between multiple objectives. [^12] The default values for $\alpha$, $\beta$ and $\gamma$ are $1.0$, $1.0$ and $1.0$, respectively. Note that we allow the area of chiplets to increase to satisfy the reach constraints. During optimization, the following constraints are considered:
@@
[…]
SA-based chiplet floorplanner. Our reach-aware floorplanner uses Sequence Pair (MurataFNK96, ) to represent spatial arrangement of chiplets in the netlist and Simulated Annealing (KirkpatrickGV83, ) to optimize the objective function. The Sequence Pair-based annealing supports five solution perturbation (move) operators with respective probabilities 0.2, 0.2, 0.2, 0.2 and 0.2:
@@
  • Op4: Reshape a chiplet. Find the chiplet with highest reach violation penalty and change its shape using the resizing algorithm in (ChenC06, ). The example of Figure 8 (top) shows how a reach violation for the net connecting chiplets E and B (red dashed arrow) is fixed by aligning the right boundary of B with the right boundary of E.
  • Op5: Expand (i.e., bloat) a chiplet. Find the chiplet with highest reach violation penalty and bloat it to occupy neighboring whitespace, with bloating in each direction determined by the length required to fix the reach violation. Figure 8 (bottom) shows how reach violations for nets connecting chiplets B-G and B-F are fixed by bloating B towards the top and right. This operator helps reduce the reach violation penalty but comes at the expense of increasing both the chiplet area and the overall package area.
To enhance the performance of simulated annealing (SA), we adopt a Go-With-the-Winners (GWTW) strategy (Aldous94, ), which enables 10 parallel SA walkers to independently explore the solution space. Periodically, a synchronization phase is performed in which: (i) the best-performing threads are identified, (ii) their solutions are cloned to repopulate the thread pool and (iii) all threads resume independent exploration. This strategy balances diversification and intensification, improving convergence without incurring significant runtime overhead. We use 10 threads in our experiments, a setting that is easily supported on standard server-class machines. [^13] This number of threads is typical for modern multi-core servers and does not impose a significant hardware burden. As discussed in Section 5.1, our chiplet floorplanner operates in two modes: “standard” and “fast”. The standard mode performs 1,000,000 perturbations with a cooling rate of 0.989, while the fast mode performs 10,000 perturbations with an initial temperature of 0.0 (greedy). In our experiments, the “standard mode” is invoked after each FM or KL pass (default: 4 passes), totaling four invocations per run. The fast mode is triggered after each FM vertex move (default: 50% of all vertices) and each KL vertex swap (default: 10% of all vertices), resulting in 230 invocations per pass in our largest testcase. These parameters were empirically selected to strike a balance between runtime and solution quality.
@@ 6. Experimental Results
ChipletPart is implemented using approximately 27K lines of C++ code, building upon implementations from (BustanyGKKPW23, ). We use OpenMP (openmp, ) for parallelization, the Eigen library (eigen_lib, ) for spectral partitioning, codes from  (metis_repo, ) to run METIS and Boost (boost, ) for high precision arithmetic computations in the cost model. Additionally, we have translated the Python-based cost model from  (Graening_CATCH, ; costmodelUCLA, ) into C++ and integrated it into our ChipletPart framework. The scripts and source code of ChipletPart, including its cost model implementation, are available in  (anonymousrepo, ). We run all experiments on a Linux server with Intel Xeon E5-2690 CPU (48 threads) and 256GB RAM.
@@ 6.1. Benchmarks and Baselines
We evaluate four categories of testcases in this work: Waferscale, MemPool, Industrial Testcase and Comparison Testcases. A common issue in partitioning research is that most testcases are too small to warrant practical chiplet partitioning; existing works often focus on very small designs, overemphasizing the impact of IO and underemphasizing yield and packaging costs. To address this, we scale up the power and area of our baseline designs to sizes more representative of real commercial systems (e.g., Intel’s Ponte Vecchio (IntelPV, )). We also scale the interconnects using Rent’s Rule, [^14] Rent’s Rule (Landman71, ) states that the number of terminals $T$ in a logic block scales as $T=kC^{p}$, where $k$ is a constant, $C$ is the number of components and $p$ is the Rent parameter. Hence, if $C$ is scaled by a factor $s$, $T$ is scaled by $s^{p}$. applying constants from (Singh1994, ) for microprocessor and memory blocks. This ensures that our scaled designs have a realistic number of IOs, making them suitable for evaluating chiplet partitioning.
Our first category of testcases (Waferscale) is derived from the waferscale graph processor in (Pal2021, ). We extract synthesized IP block areas and a block-level netlist at 45 nm, then apply area and power scaling to favor partitioning. The second category (MemPool) is similarly extracted from synthesized blocks in (Cavalcante2021, ) and likewise scaled from its 45 nm block-level netlist. The third category (Industry Testcase) is based on a 16 nm design provided by a semiconductor company, which we also scale to a suitable size. Finally, the Comparison Testcases based on publicly available information for large commercial designs come from (Li24, ). The original implementations were in various nodes and these designs are sufficiently large that they do not require further scaling.
@@
Table 4 shows the specific testcase configurations that we use; see also the design-specific discussion below. In the table, the WS testcases are based on the Waferscale design, the MP testcase is based on the MemPool design, the TC testcase is based on the Industry design and EPYC and GA100 are the Comparison Testcases. The specific parameters are explained as follows.
Waferscale Testcases. The waferscale design is organized into a grid of tiles connected in a 2D mesh configuration with neighbor-to-neighbor communication between tiles. Each tile comprises 48 IP blocks. Those blocks include a router for tile-to-tile connections, a crossbar for intra-tile connections, 4 large shared memory blocks connected to the crossbar, and 14 cores each with associated bus and private memory blocks. Figure 9 gives a system block diagram. We have implemented a testcase generator (see (anonymousrepo, )) which allows us to choose the number of tiles in a given testcase, along with the number of cores and memories per tile. Area and power scaling are also provided via the testcase generator. As noted, Table 4 shows the waferscale configurations used in our case studies.
@@
[…]
MemPool Testcase. The full MemPool design consists of 4 MemPool groups, each containing 16 MemPool tiles. These tiles each consist of cores and memory along with supporting logic. Note that while each tile contains 4 cores and 16 memory banks, we do not split tiles for partitioning. For our MemPool testcase, we use a single MemPool group, i.e., taken at just above the tile level of the hierarchy. Based on 16 tiles and 24 blocks related to (remote, local, AXI etc.) interconnect, our block-level netlist for MemPool group has 40 IP blocks to use in partitioning. A diagram is available in (anonymousrepo, ); see also Figure 3(a) in (Cavalcante2021, ).

### v3

@@ 33  Apply mutation to each offspring $\omega^{o}_{x}$ with mutation probability $p_{m}$;
34  Construct the next generation by selecting the top $\sigma$ elite genomes from the current generation, along with $k_{pop}$ offspring;
35
@@ 4. Chiplet Partitioning Drivers
Smaller chiplets can potentially bring lower costs and improved yield, while technology heterogeneity offers better power and performance. The overhead of inter-chiplet I/O also affects how a system should be partitioned into chiplets. In this section, we briefly discuss these factors.
@@ 4.1. Chiplet Cost Model
We use the open-source cost model from (Graening_CATCH, ; costmodelUCLA, ). It computes cost and yield based on the chiplet/interposer area (dependent on the technology node), assembly processes (dependent on bonding parameters), and I/O placement constraints (dictated by netlist connectivity and reach). The model calculates the individual cost and yield of each chiplet along with the interposer, and then aggregates them for the full assembly cost. We consider designs to be high volume in our studies, to minimize the impact of NRE. The cost model is summarized by Equation 1. For further details, see (Graening_CATCH, ). Benefits of using the cost model in (Graening_CATCH, ) are discussed in Section 6.2 with results in Tables 18 and 6.
@@ 4.2. Chiplet Power Model
We assume that performance of chiplets is preserved across technology nodes but that the power changes. We further assume that inter-chiplet communication latency is small enough to not influence the architecture. If certain inter-block interfaces are highly latency-sensitive, then those blocks should be merged in our framework to map them to the same chiplet. More complex performance models as in (ChenLZZL23, ) are possible but not explored in this work.
Power is calculated as the sum of block power and I/O power. If all blocks are in the same chiplet, there will be no additional power due to I/O. However, if the blocks are placed into multiple partitions (chiplets), some additional power will be consumed by the added I/O cells. We scale the block power according to the scaling factors in (Stilmaker2017, ) for different technology nodes.
4.3. I/O Reach Model
We use an I/O cell model where each I/O cell has both an area and a “reach” which is the maximum wire length that can be driven by the I/O cell. The reach depends on the I/O transceiver design and is often dictated by the inter-chiplet I/O standard. For example, Universal Chiplet Interconnect Express (UCIe) (UCIe1p0, ) guarantees a reach of 2 mm for advanced packages while standard packages (such as organic substrates) support larger reaches of up to 25 mm.
Since we assume a fixed system architecture, our partitioner does not modify the network structure or insert pipeline stages to compensate for long inter-chiplet wires. We also assume a standardized I/O cell with a fixed maximum reach instead of more complicated I/O structures. These assumptions help keep the partitioning problem well-scoped and tractable. We leave sophisticated I/O planning including pass-through connections, inserting buffers on the substrate, multiple I/O types per design, etc. to future work. However, users can explore multiple architectural variants and I/O types for the same design. For example, in our industry testcase (Section 6.1), we evaluate two versions that differ only in their crossbar configurations — and in Figure 14, we examine the impact of different I/O cell types on partitioning.
Refer to caption An image showing squares in a 3x3 configuration representing chiplets with connections shown as colored lines of various lengths to indicate the different reach requirements fore connections from one chiplet to another in the grid.
@@ 4.3. I/O Reach Model
Consider the example in Figure 4, and assume that a bundle of wires is connected from the side of one chiplet to the side of another. If the chiplets are 5 mm on a side and spaced apart by 1 mm, then the short black connections require a reach of at least $1~\text{mm}$, while the orange, blue, and green connections require reaches of at least 7 mm, 13 mm, and 19 mm respectively. Small values of reach constrain the floorplan, while larger values of reach come at the cost of more expensive I/O cells.
I/Os are discretized depending on the standard. For instance, a single UCIe I/O cell is as large as 0.88 $\text{mm}^{2}$ (UCIe1p0, ) while for parallel signaling (e.g., Advanced Interface Bus (AIB) or (Pal2021, )), an I/O cell can be as small as 157 $\text{$\mu$m}^{2}$ (PalIO2021, ). This discretization can add complexity for chiplet partitioning, as the I/O (i.e., net cut) cost calculation follows a stepwise curve as a function of net cut.
@@ 5. Chiplet Partitioning Core
In this section, we first discuss our chiplet-cost-driven partitioning approach Core-ChipletPart in Section 5.1. Then, we present our reach-aware chiplet floorplanning approach in Section 5.2.
@@ 5.1. Cost-driven Partitioning: Core-ChipletPart
Our Core-ChipletPart does not adopt the widely-used multilevel approach since there are typically only a few hundreds of IP blocks in a block-level netlist. As shown in Figure 5, we (i) compute a large pool of initial partitioning solutions using a variety of graph partitioning algorithms; (ii) prune poor-quality partitioning solutions from the pool; and (iii) perform floorplan-aware Fiduccia–Mattheyses (1585498, ) (FM)-based and Kernighan-Lin (HELSGAUN2000106, ) (KL)-based refinement. We then output the solution with the best cost.
Refer to caption A flowchart showing the Core-ChipletPart partitioning flow.
Figure 5. Core-ChipletPart partitioning flow. We use multiple techniques to generate the initial partitions, then we prune out the worst-performing initializations before running refinement.
Generation of a partitioning pool. Similar to the initial partitioning stage in widely-used multilevel partitioning approaches (KarypisAKS99, ; BustanyGKKPW23, ), we compute a pool of initial partitioning solutions so as to explore a larger solution space. [^9] However, unlike multilevel approaches (KarypisAKS99, ; BustanyGKKPW23, ), Core-ChipletPart omits the coarsening stage due to the relatively small number of vertices (i.e., 384 IP blocks in our largest testcase, as shown in Table 4) in the netlist graph. Specifically, we use the following graph partitioning methods.
  • Spectral partitioning: Because the block-level netlist is both small and sparse, we apply spectral partitioning (von2007tutorial, ). We first compute the spectral embedding using the two smallest nontrivial eigenvectors, then cluster the IP blocks with the K-means algorithm (likas2003global, ). [^10] In our implementation, we use a parallel K-means method with smart initialization (K-means++ (bahmani2012scalable, )), and we set $K=4$ by default.
  • High-degree nodes expansion: Our node expansion approach follows (zhang2017graph, ). In the block-level netlist, high-degree nodes (e.g., crossbars) are distributed across different chiplets, and then expanded via breadth-first search. If multiple chiplets qualify for expansion, the block is assigned to the chiplet with which it has the strongest connection.
@@
Partitioning refinement. We use two strategies for refinement: (i) FM and (ii) KL. Both of these processes are shown in Figure 6. Each potential FM move or KL swap triggers a call to the chiplet cost model to obtain the cost (gain) of an updated partitioning solution. This evaluation includes generating a floorplan solution using the “fast” mode of our SA-based reach-aware chiplet floorplanner, followed by calculating the corresponding cost. After completing each FM or KL pass, the “standard” mode of the SA-based reach-aware chiplet floorplanner is employed to search for an improved floorplan solution for the subsequent pass. Further details about our SA-based reach-aware chiplet floorplanner are provided in Section 5.2. We adopt the K-way FM implementation from TritonPart (BustanyGKKPW23, ), while our KL refinement (klrefine, ) follows the methodology described in (HELSGAUN2000106, ). In our implementation, we apply KL after FM to further improve solution quality, as KL explores a broader neighborhood by enabling pairwise vertex swaps — compared to FM’s single-vertex moves.
Refer to caption An flowchart showing the behavior of partition refinement.
Figure 6. Floorplan-aware FM and KL-based refinement. In our implementation, we run FM first, followed by KL.
@@ 5.2. Reach-aware Chiplet Floorplanning
Chiplet floorplanning is a crucial step, as it determines the location and shape of each chiplet on the interposer. In contrast to the well-studied macro placement, chiplet floorplanning must also consider I/O-feasibility. More specifically, the wirelength of a net must not exceed the reach specified by the corresponding I/O cell (Section 4). Figure 7 illustrates the wirelength (measured in terms of Manhattan distance) calculation for net $e$ connecting chiplets $A$ and $B$. Assuming that the bitwidth of net $e$ is $n_{e}$ and the area of I/O cells is $A_{IO}$, then the wirelength of net $e$ ($length(e)$) in the figure is calculated as: (i) $l=\sqrt{w^{2}+2A_{IO}}-w$ and (ii) $length(e)=h+2l$, where $l$ is the depth needed for all the I/O cells. [^12] The detailed code for more general cases is available in (reach, ). Then the reach violation penalty for net $e$ is $n_{e}\times max(length(e)-reach(e),0.0)$. The reach violation penalty for a chiplet is defined as the summation of the reach violation penalties for all the nets connected to the chiplet.
Refer to caption Diagram showing two chiplets (A and B) represented by rectangles placed next to each other with distances shown in the corresponding equation marked.
Figure 7. Illustration of wirelength computation ($length(e)$). The yellow region indicates the placement area for I/O cells that will satisfy the reach constraint.
@@
where $WL_{reach}$, $A_{C}$, and $A_{P}$ respectively denote the reach violation penalty, the area of chiplets, and the area of the package. $\alpha$, $\beta$ and $\gamma$ are user-defined coefficients that can be modified to achieve a desired trade-off between the different objectives in Equation 3.131313The default values for $\alpha$, $\beta$ and $\gamma$ are $1.0$, $1.0$ and $1.0$, respectively. Note that we allow the area of chiplets to increase to satisfy the reach constraints. During optimization, the following constraints are considered:
@@
Image showing different possibilities for resizing chiplets.
@@
SA-based chiplet floorplanner. Our reach-aware floorplanner uses Sequence Pairss (MurataFNK96, ) to represent a spatial arrangement of chiplets in the netlist and Simulated Annealing (KirkpatrickGV83, ) to optimize the objective function (Equation 3). Our annealer supports five solution perturbation (move) operators, each selected with equal probability (0.2).
@@
  • Op4: Reshape a chiplet. Identify the chiplet with the highest reach-violation penalty and modify its shape using the resizing algorithm described in (ChenC06, ). Figure 8 (top) illustrates how a reach violation for the net connecting chiplets E and B (red dashed arrow) is resolved by aligning the right boundary of B with that of E.
  • Op5: Expand (i.e., bloat) a chiplet. Identify the chiplet with the highest reach-violation penalty and expand it into neighboring whitespace, where the expansion in each direction is determined by the extent required to eliminate the violation. Figure 8 (bottom) shows how reach violations for nets connecting chiplets B–G and B–F are resolved by expanding B toward the top and right. This operator reduces reach-violation penalties at the cost of increasing the chiplet area and potentially the overall package area as well.
To enhance the performance of simulated annealing (SA), we adopt a Go-With-the-Winners (GWTW) strategy (Aldous94, ), which enables 10 parallel SA walkers to independently explore the solution space. Periodically, a synchronization phase is performed in which: (i) the best-performing threads are identified, (ii) their solutions are cloned to repopulate the thread pool and (iii) all threads resume independent exploration. This strategy balances diversification and intensification, improving convergence without incurring significant runtime overhead. We use 10 threads in our experiments, a setting that is easily supported on standard server-class machines. [^14] This number of threads is typical for modern multi-core servers and does not impose a significant hardware burden. As discussed in Section 5.1, our chiplet floorplanner operates in two modes: “standard” and “fast”. The standard mode performs 1,000,000 perturbations with a cooling rate of 0.989, while the fast mode performs 10,000 perturbations with an initial temperature of 0.0 (greedy). In our experiments, the “standard mode” is invoked after each FM or KL pass (default: 4 passes), totaling four invocations per run. The fast mode is triggered after each FM vertex move (default: 50% of all vertices) and each KL vertex swap (default: 10% of all vertices), resulting in 230 invocations per pass in our largest testcase. These parameters are empirically selected to maintain a balance between runtime and solution quality.
@@ 6. Experimental Results
ChipletPart is implemented using approximately 27K lines of C++ code, building upon implementations from (BustanyGKKPW23, ). We use OpenMP (openmp, ) for parallelization, the Eigen library (eigen_lib, ) for spectral partitioning, codes from  (metis_repo, ) to run METIS, and Boost (boost, ) for high precision arithmetic computations in the cost model. Additionally, we have translated the Python-based cost model from  (Graening_CATCH, ; costmodelUCLA, ) into C++ and integrated it into our ChipletPart framework. The scripts and source code of ChipletPart, including its cost model implementation, are available in  (anonymousrepo, ). We run all experiments on a Linux server with Intel Xeon E5-2690 CPU (48 threads) and 256GB RAM.
@@ 6.1. Benchmarks and Baselines
We evaluate four categories of testcases in this work: Waferscale, MemPool, Industrial Testcase, and Comparison Testcases. A common issue in chiplet partitioning research is that most testcases are too small to warrant practical chiplet partitioning; existing works often focus on very small designs, overemphasizing the impact of I/O and underemphasizing yield and packaging costs. To address this, we scale up the power and area of our baseline designs to sizes more representative of real commercial systems (e.g., Intel’s Ponte Vecchio (IntelPV, )). We also scale the interconnects using Rent’s Rule, [^15] Rent’s Rule (Landman71, ) states that the number of terminals $T$ in a logic block scales as $T=kC^{p}$, where $k$ is a constant, $C$ is the number of components and $p$ is the Rent parameter. Hence, if $C$ is scaled by a factor $s$, $T$ is scaled by $s^{p}$. applying constants from (Singh1994, ) for microprocessor and memory blocks. This ensures that our scaled designs have a realistic number of I/Os, making them suitable for evaluating chiplet partitioning.
Our first category of testcases (Waferscale) is derived from the waferscale graph processor in (Pal2021, ). We extract synthesized IP block areas and a block-level netlist at 45 nm, then apply area and power scaling to favor partitioning. The second category (MemPool) is similarly extracted from synthesized blocks in (Cavalcante2021, ) and likewise scaled from its $45\text{nm}$ block-level netlist. The third category (Industry Testcase) is based on a $16\text{nm}$ design provided by a semiconductor company, which we also scale to a suitable size. Finally, the Comparison Testcases are based on publicly available information for large commercial designs, and come from (Li24, ). The original implementations span multiple technology nodes, and these designs are sufficiently large that no additional scaling is required.
@@
Table 4 shows the specific testcase configurations that we use; see also the design-specific discussion below. In the table, the WS testcases are based on the Waferscale design, the MP testcase is based on the MemPool design, the TC testcase is based on the Industry design, and EPYC and GA100 are the Comparison Testcases. The specific design configurations are explained as follows.
Waferscale Testcases. The Waferscale design is organized into a grid of tiles connected in a 2D mesh configuration with neighbor-to-neighbor communication between tiles. Each tile comprises 48 IP blocks. Those blocks include a router for tile-to-tile connections, a crossbar for intra-tile connections, four large shared memory blocks connected to the crossbar, and 14 cores each with associated bus and private memory blocks. Figure 9 gives a system block diagram. We have implemented a testcase generator (see (anonymousrepo, )) which allows us to choose the number of tiles in a given testcase, along with the number of cores and memories per tile. Area and power scaling are also provided via the testcase generator. As noted, Table 4 shows the Waferscale configurations used in our case studies.
@@
Refer to caption Image shows a block diagram of the waferscale testcase used in the paper. The image shows a tiled configuration where the blocks within each tile are arranged around a centralized crossbar.
@@
MemPool Testcase. The full MemPool design consists of four MemPool groups, each containing 16 MemPool tiles. These tiles each consist of cores and memory along with supporting logic. Note that while each tile contains four cores and 16 memory banks, we do not split tiles for partitioning. For our MemPool testcase, we use a single MemPool group, i.e., taken at just above the tile level of the hierarchy. Based on 16 tiles and 24 blocks related to (remote, local, AXI etc.) interconnect, our block-level netlist for MemPool group has 40 IP blocks to use in partitioning. A diagram is available in (anonymousrepo, ); see also Figure 3(a) in (Cavalcante2021, ).
@@
Refer to caption Image showing the blocks in the AMD EPYC 7282 testcase. Blocks are represented by rectangles and connections are shown with colored lines.